OpenStudy (cassieforlife5):

Find the limit as x approaches 0: 2sin^3 3x/-x^3 Any help would be appreciated! Thank you!

ganeshie8 (ganeshie8):

Hint : \[ \dfrac{\sin^3(3x)}{x^3} = 27 \cdot \dfrac{\sin^3(3x)}{(3x)^3}\]

OpenStudy (gorv):

\[\lim_{x \rightarrow 0}\frac{ sinx }{ x }=1\]

OpenStudy (gorv):

does it ring a bell

OpenStudy (cwrw238):

so can we say that limit of sin^3 3x / / 3x^3 has a limit of 1 as x approaches 0?

OpenStudy (cwrw238):

if so then the required limit is -2 8 27 = -54

OpenStudy (cwrw238):

* -2 * 27 = -54

OpenStudy (cwrw238):

i think this is valid - limits are not my strong point

OpenStudy (cassieforlife5):

So I'm going with -54 on that question. Could you also check my answer? I'm supposed to find the derivative of f(x)= ((3x^2)/2) -x I got 3x-1 for my answer

ganeshie8 (ganeshie8):

\(\LARGE \color{Red}{\checkmark}\)

OpenStudy (cassieforlife5):

Thanks to everyone so much!